# Feature Request Analysis and Processing: Tasks, Techniques, and Trends

### v1

@@ 5.1. Feature Requests Identification
Feature requests are often embedded in diverse communication channels and platforms, including issue tracking systems, support forums, app store reviews, and social media comments. Figure 6 illustrates the various forms of primary sources for feature requests as identified in existing research. As shown, feature requests originate from a wide range of channels, including app reviews, issue reports, user feedback, Twitter posts, development emails, customer suggestions, and chat messages. Some sources explicitly present requests in the form of feature requests, change requests, or enhancement reports. Among these, app reviews (61) and issue reports (17) are the most frequently used sources for extracting feature requests. Although user feedback varies in format across different platforms, it typically contains feature requests either explicitly or implicitly. For example, Iacob and Harrison (Iacob and Harrison, 2013) found that approximately 23.3% of application reviews include feature requests—often in the form of user suggestions for new functionality or improvements to existing features. Similarly, Williams and Mahmoud (Williams and Mahmoud, 2017) reported that around 51% of user comments on Twitter provide useful information, comprising 27% bug reports and 24% user requests. In addition to feature requests, user feedback may also include bug reports and insights into user experience.
Feature request identification refers to the task of identifying and extracting descriptions of desired new software functionalities from such heterogeneous and often unstructured data. Given the sheer scale of user feedback—especially in commercial platforms such as the App Store, which hosted over 1.91 million apps, with an average of 1,342 new apps released every day and generated an immense volume of user reviews by early 2025  [^4] https://www.tekrevol.com/blogs/apple-app-store-statistics/?utm_source=chatgpt.com—manual extraction is infeasible for real-world applications, despite its potential accuracy.
@@
Table 4. Classification of Approaches for Feature Requests Extraction
@@
Unsupervised Rule Matching (Carreño and Winbladh, 2013)(Iacob and Harrison, 2013)
@@
[…]
@@ 5.1.2. Semi-supervised Learning-based Approaches
Semi-supervised methods for feature request extraction typically encompass techniques such as self-training (Yarowsky, 1995), Rasco (Wang et al., 2008), Rel-Rasco (Yaslan and Cataltepe, 2010), and active learning (Bowring et al., 2004). For example, Deocadez et al. (Deocadez et al., 2017) employed three semi-supervised algorithms—self-training, Rasco, and Rel-Rasco—to categorize app reviews into bugs, requests, and other categories. They evaluated these methods using four base classifiers: KNN, C4.5, SMO, and Naive Bayes. Their results demonstrated that semi-supervised approaches can substantially reduce the need for labeled data, achieving comparable classification accuracy using only 30% of the labeled data required by fully supervised methods. Similarly, Dhinakaran et al. (Dhinakaran et al., 2018) utilized active learning to extract feature requests from user reviews, applying three uncertainty sampling strategies—lowest confidence, margin sampling, and entropy—to identify the most informative samples for annotation. Their study found that active learning notably decreases labeling effort while maintaining performance on par with traditional supervised learning.
@@ 5.1.3. Supervised Learning-based Approaches
The most widely adopted approach is supervised learning-based classification, which can be divided into two categories based on the granularity: sentence-level and text-level classification, as summarized in Table 4. Text-level classification treats each user feedback entry as a whole and assigns labels such as bug reports, feature requests, praise, user experience, inquiries, and others (Abbas et al., 2024; Al-Hawari et al., 2021; Arya et al., 2019). This approach effectively identifies feature requests and filters out irrelevant reviews. Sentence-level classification breaks down feedback into individual sentences and predicts labels like information provision, information search, feature request, problem detection, explanation, solution, and others for each sentence. This fine-grained approach helps localize feature requests more precisely within the feedback and further excludes non-relevant information.
Figure 7. The common text classification process (“Feature Extraction” here means the process of transforming natural language into machine-interpretable features for learning tasks.).
The typical workflow for text classification is illustrated in Figure 7. Large volumes of user feedback—such as app reviews and issue reports from open-source platforms—are preprocessed using natural language processing (NLP) techniques including tokenization, punctuation removal, stop-word elimination, lemmatization, and stemming. The processed text is then vectorized using methods like bag-of-words (BOW) (e.g., TF-IDF, term frequency), n-grams, or word embeddings (e.g., Word2Vec (Mikolov et al., 2013), GloVe (Pennington et al., 2014)).
Beyond textual features, many studies incorporate auxiliary metadata such as review topics, rating levels, text length, sentence position, app size, installation counts, and reviewer numbers (Guzman et al., 2015; Maalej and Nabil, 2015; Peng et al., 2016; Arya et al., 2019; Aslam et al., 2020; Merten et al., 2016a; Tizard et al., 2019). Additional auxiliary features mined via NLP include keywords (Maalej and Nabil, 2015; Peng et al., 2016; Song et al., 2020; Merten et al., 2016a), sentiment analysis (Williams and Mahmoud, 2017; Guzman et al., 2015; Maalej and Nabil, 2015; Peng et al., 2016; Aslam et al., 2020; Yadav et al., 2020; Panichella et al., 2015, 2016), semantic rules (Maalej and Nabil, 2015; Song et al., 2020), heuristic sentence identification (Panichella et al., 2015, 2016; Sorbo et al., 2015), lexical features (Panichella et al., 2016), and part-of-speech tagging (Panichella et al., 2015, 2016; Sorbo et al., 2015).
Finally, classifiers based on traditional machine learning or deep learning models are trained to predict categories of feedback  [^5] We will introduce LLM-based approaches separately in a later section.. Due to the often imbalanced distribution of categories in user reviews, some studies have validated resampling techniques to balance datasets and improve classifier performance (Arya et al., 2019; Khan et al., 2020; Tizard et al., 2019).
Figure 8. The common text classification algorithms
@@
Early approaches employed traditional classifiers such as naïve Bayes (NB), multilayer naïve Bayes (MNB), support vector machines (SVM), logistic regression (LR), decision tree (DT), random forest (RF), and maximum entropy (MaxEnt), often combined with BOW-based vectorization like TF-IDF, TF, and n-grams. With the rise of deep learning, models such as multilayer perceptron (MLP), convolutional neural networks (CNN) (LeCun et al., 1998), text recurrent neural networks (TextRNN) (Lai et al., 2015), long short-term memory networks (LSTM) (Hochreiter and Schmidhuber, 1997), and fastText (Grave et al., 2018) have been increasingly applied. These are often used in conjunction with word embedding methods, such as Word2Vec, GloVe, and BERT (Devlin et al., 2018). Figure 8 shows the yearly distribution of classification algorithms in literature, with the x-axis representing algorithms, the y-axis representing years, and circle numbers denoting citation counts. Traditional classifiers, such as NB and DT, remain the most common, followed by RF and SVM. Deep learning methods began emerging around 2019, with CNNs becoming the most prevalent deep model for feature request extraction. Performance varies depending on conditions and datasets. Guzman et al. (Guzman et al., 2015) showed that ensemble classifiers outperform single classifiers. Stanik et al. (Stanik et al., 2019) observed that traditional classifiers excel on small datasets, whereas deep learning models perform better with large datasets. Khan et al. (Khan et al., 2024) evaluated 14 algorithms, finding that multi-layer perception (MLP) and RF achieve the highest average accuracy. In contrast, deep models like CNN, BiGRU, and GRU better capture complex textual patterns, excelling in precision, recall, and F-score for certain tasks. Ali Khan et al. (Khan et al., 2020) cautioned that deep learning may overfit on small samples, emphasizing appropriate classifier choice. To address imbalanced classes, Young et al. (Young et al., 2020) investigated resampling techniques (random undersampling, SMOTE oversampling, random oversampling, and SMOTE-Tomek) on five classifiers, finding that oversampling often yields better results. Nafees et al. (Nafees and Rehman, 2021) integrated semantic textual features and metadata into a gradient boosting machine (GBM) model with SMOTE-Tomek to predict approval probabilities for feature requests. Abbas et al. (Abbas et al., 2024) combined RF with an accuracy sliding window (ASW) technique, improving classification accuracy while reducing tree counts, providing robust performance and new insights.
Many studies enhance classification by incorporating extra features extracted via NLP techniques, such as keywords, sentiment analysis, tense, semantic rules, and part-of-speech tagging. Keyword-based features have been particularly effective, as bug reports and feature requests exhibit distinct keyword patterns  (Maalej and Nabil, 2015; Peng et al., 2016; Merten et al., 2016a). Sentiment analysis, using tools like SentiStrength (Thelwall et al., 2012), SnowNLP, or Senti4SD (Claytor and Servant, 2018), classifies reviews as positive, negative, or neutral. Peng et al. (Peng et al., 2016) demonstrated that sentiment classification improves feature request identification in app reviews. However, Williams and Mahmoud (Williams and Mahmoud, 2017) found sentiment less effective in classifying Twitter feedback, possibly due to its narrative style lacking clear emotional signals, unlike app store reviews which often contain complaints. Al-Hawari et al. (Al-Hawari et al., 2021) proposed combining association rule mining and clustering to extract high-frequency functional terms (e.g., multi-device synchronization, dark mode) as feature labels and clustered feature requests into maintenance tasks like “functional enhancement” and “user experience optimization,” focusing on semantic parsing from an implementation perspective rather than traditional sentiment analysis.
@@ 5.1.4. Pre-trained Language Model Approaches
Despite their effectiveness in handling large-scale data, PLMs exhibit a black-box nature, making their decision processes opaque. This opacity hinders developers’ ability to fully trust and utilize model outputs, especially considering potential biases in the training data. To address these issues, recent research has emphasized the importance of interpretability and transparency. For instance, Gambo et al. (Gambo et al., 2024) proposed a method that combines a BiLSTM with attention mechanisms and leverages local interpretable model-agnostic explanations (LIME) and SHapley Additive exPlanations (SHAP) to improve feature request identification. This approach outperforms SVM and CNN in accuracy and F1 score. Additionally, Biswas et al. (Biswas et al., 2024) developed the IARCT method, which quantifies the contribution of each word via Shapley values and provides contextual explanations for Transformer-based classification decisions. However, challenges remain due to computational costs, subjective interpretability, limited generalization, and small dataset sizes.
@@ 5.1.5. General Large Language Model Approaches
Ghosh et al. (Ghosh et al., 2024) evaluated the performance of three LLMs—BERT, DistilBERT, and GEMMA—on 3,200 Google Play Store user reviews. Their study confirmed the effectiveness of LLMs in parsing unstructured text and accurately identifying feature requests, showcasing the potential of LLMs to transform user feedback into actionable requirements. Building on this, Assi et al. (Assi et al., 2024) introduced the LLM-Cure framework, which derives feature improvement suggestions from user reviews via a two-phase approach. First, it extracts high-frequency functional features by semantically clustering batch reviews and normalizing cross-app features using word embeddings and cosine similarity. Then, for low-rated reviews of a target app, it retrieves corresponding high-rated competitor reviews to create contextual input for generating specific improvement suggestions through retrieval-augmented generation. This strategy enhances the relevance and specificity of feature requests by incorporating competitive insights. Addressing challenges such as labor-intensive manual analysis and low clustering accuracy, Ren et al. (Ren et al., 2024) proposed a hybrid method that combines Latent Dirichlet Allocation (LDA) topic modeling for user review classification with a GPT-4–based generation approach for categories with sparse reviews. By incorporating example-based prompting and structuring generated requirements as hierarchical trees, their method improves the systematic identification of feature requests, particularly in datasets with limited review volume. Furthermore, Wei et al. (Wei et al., 2024) conducted a comparative study of feature refinement approaches based on app store data and LLMs. Selecting 20 high-level root features (half existing and half novel) across diverse domains, they utilized App Store search and GPT-4 to generate two-level feature trees. Through independent evaluation by three authors on relevance and feasibility, results indicated that LLMs excel at discovering novel, app-specific features but may struggle with feasibility, whereas app store–based methods perform better in cross-domain feature recommendations and feasibility validation. Their integrated approach, implemented in the piStar(Pimentel and Castro, 2018) tool, reduces effort and inspires creativity, providing practical guidance for selecting appropriate feature request identification methods.
@@ 5.1.6. Other Approaches
Rahimi et al. (Rahimi and Cleland-Huang, 2014) proposed a semi-automated method for identifying feature requests from online forums by clustering user comments via incremental diffusion clustering and applying Apriori association rule mining to identify stakeholder-relevant topic clusters. These clusters reveal quality issues related to performance, security, and usability, and help create personas that guide architectural design and reduce costly late-stage refactoring. Zhang et al. (Zhang et al., 2019) developed a hybrid similarity-based approach to classify issue reports as bugs, feature requests, or others by combining normalized text similarities between unlabeled issue reports, user reviews, and labeled reports with weighted contributions. Shi et al. (Shi et al., 2020) introduced FRMiner, which uses a deep Siamese network to determine whether chat messages belong to the same class by learning their similarity, enabling feature requests identification. Panichella et al. (Panichella et al., 2015) combined manual rule matching of 246 linguistic patterns with sentiment analysis and J48 classification to categorize user reviews into information provision, search, feature requests, problem detection, and others. Gribkov et al. (Gribkov and Yekhlakov, 2020) presented the Trans-LSTM model to extract (topic, description) expressions from user review sentences, classifying them into bug reports, feature requests, and positive or negative features. Their use of Trans-CNN and Trans-LSTM outperformed prior models. Shi et al. (Shi et al., 2021) proposed an automated method for identifying feature requests from development emails using semantic sequence patterns generated from fuzzy rules and email context represented via a 2-gram model, validated on Ubuntu community emails. Izadi et al. (Izadi et al., 2022) introduced a multi-task approach that predicts both issue type (bug, feature, question) using fine-tuned BERT, compared with RF and SVM, and issue priority (high, medium, low), incorporating the predicted issue type to enhance cross-task learning.
@@ 5.2. Quality Review
Refer to caption(Tudor and Walter, 2006; Avesani et al., 2005; Sillitti and Succi, 2005)(Tudor and Walter, 2006; Avesani et al., 2005)(Chen et al., 2014; Keertipati et al., 2016; Morales-Ramirez et al., 2017; Etaiwi et al., 2020)(Aslam et al., 2020)(Laurent and Cleland-Huang, 2009)(Tudor and Walter, 2006; Avesani et al., 2005)
Figure 9. Feature Request Quality Model
Existing research on feature request review primarily centers on defining quality standards (Heck and Zaidman, 2015; Shi et al., 2016) and employing these criteria to assess the quality of requests (Heck and Zaidman, 2017, 2013). In addition, a growing body of work has proposed automated techniques to detect duplicate (Aversano, 2019; Gu et al., 2011; Li et al., 2017; Heck and Zaidman, 2014), redundant (Shi et al., 2016), and ambiguous (Aslam et al., 2020) feature requests, aiming to streamline the review process and improve the overall quality of requirements. We synthesize the quality standards identified in existing studies, as illustrated in Figure 9.

### v2

@@ 5.1. Feature Requests Identification
Feature requests are often embedded in diverse communication channels and platforms, including issue tracking systems, support forums, app store reviews, and social media comments. Figure 7 illustrates the various forms of primary sources for feature requests as identified in existing research. As shown, app reviews (68) and issue reports (19) are the most frequently used sources for extracting feature requests. Although user feedback varies in format across different platforms, it typically contains feature requests either explicitly or implicitly. For example, Iacob and Harrison (Iacob and Harrison, 2013) found that approximately 23.3% of application reviews include feature requests—often in the form of user suggestions for new functionality or improvements to existing features. Similarly, Williams and Mahmoud (Williams and Mahmoud, 2017) reported that around 51% of user comments on Twitter provide useful information, comprising 27% bug reports and 24% user requests. In addition to feature requests, user feedback may also include bug reports and insights into user experience.
Feature request identification refers to the task of identifying and extracting descriptions of desired new software functionalities from such heterogeneous and often unstructured data. Given the sheer scale of user feedback—especially in commercial platforms such as the App Store, which hosted over 1.91 million apps, with an average of 1,342 new apps released every day and generated an immense volume of user reviews by early 2025  [^4] https://www.tekrevol.com/blogs/apple-app-store-statistics—manual extraction is infeasible for real-world applications, despite its potential accuracy.
@@
Table 3. Classification of Approaches for Feature Requests Extraction.
@@
Rule-based (Carreño and Winbladh, 2013)(Iacob and Harrison, 2013)
@@
(Abbas et al., 2024) (Al-Hawari et al., 2021) (Arya et al., 2019) (Aslam et al., 2020) (Dąbrowski et al., 2019) (Guzman et al., 2015) (Izadi et al., 2022) (Jha and Mahmoud, 2017) (Kallis et al., 2019) (Kallis et al., 2021) (Khan, 2019) (Khan et al., 2020) (Khan et al., 2024) (Maalej et al., 2016) (Maalej and Nabil, 2015) (McIlroy et al., 2016) (Nafees and Rehman, 2021) (Nayebi et al., 2018) (Peng et al., 2016) (Shi et al., 2020) (Song et al., 2020) (Stanik et al., 2019) (Tizard, 2019) (tong et al., 2018) (Villarroel et al., 2016) (Williams and Mahmoud, 2017) (Yadav et al., 2020) (Young et al., 2020)
@@ 5.1.2. Semi-supervised Learning-based Approaches
Semi-supervised methods for feature request extraction typically encompass techniques such as self-training (Yarowsky, 1995), Rasco (Wang et al., 2008), Rel-Rasco (Yaslan and Cataltepe, 2010), and active learning (Bowring et al., 2004). For example, Deocadez et al. (Deocadez et al., 2017) employed three semi-supervised algorithms—self-training, Rasco, and Rel-Rasco—to categorize app reviews into bugs, requests, and other categories. They evaluated these methods using four base classifiers: K-Nearest Neighbors (KNN), C4.5, SMO, and Naïve Bayes (NB). Their results demonstrated that semi-supervised approaches can substantially reduce the need for labeled data, achieving comparable classification accuracy using only 30% of the labeled data required by fully supervised methods. Similarly, Dhinakaran et al. (Dhinakaran et al., 2018) utilized active learning to extract feature requests from user reviews, applying three uncertainty sampling strategies—lowest confidence, margin sampling, and entropy—to identify the most informative samples for annotation. Their study found that active learning notably decreases labeling effort while maintaining performance on par with traditional supervised learning.
@@ 5.1.3. Supervised Learning-based Approaches
Supervised learning remains the most widely used approach for identifying feature requests from user feedback. Depending on granularity, methods operate at either the text or sentence level, labeling entries or individual sentences as feature requests, bug reports, or other categories (Abbas et al., 2024; Al-Hawari et al., 2021; Arya et al., 2019).
Figure 8. The Common Text Classification Algorithms.
@@
As shown in Figure 8, traditional classifiers, such as NB and Decision Tree (DT), remain the most common, followed by Random Forest (RF) and Support Vector Machines (SVM). Deep learning methods began emerging around 2019, with Convolutional Neural Networks (CNN) becoming the most prevalent deep model for feature request extraction. Performance varies depending on conditions and datasets. Guzman et al. (Guzman et al., 2015) showed that ensemble classifiers outperform single classifiers. Stanik et al. (Stanik et al., 2019) observed that traditional classifiers excel on small datasets, whereas deep learning models perform better with large datasets. Khan et al. (Khan et al., 2024) evaluated 14 algorithms, finding that MLP and RF achieve the highest average accuracy. In contrast, deep models like CNN, BiGRU, and GRU better capture complex textual patterns, excelling in precision, recall, and F-score for certain tasks. Ali Khan et al. (Khan et al., 2020) cautioned that deep learning may overfit on small samples, emphasizing appropriate classifier choice. To address imbalanced classes, Young et al. (Young et al., 2020) investigated resampling techniques (random undersampling, SMOTE oversampling, random oversampling, and SMOTE-Tomek) on five classifiers, finding that oversampling often yields better results. Nafees et al. (Nafees and Rehman, 2021) integrated semantic textual features and metadata into a gradient boosting machine (GBM) model with SMOTE-Tomek to predict approval probabilities for feature requests. Abbas et al. (Abbas et al., 2024) combined RF with an accuracy sliding window (ASW) technique, improving classification accuracy while reducing tree counts, providing robust performance and new insights.
Beyond textual features, many studies incorporate auxiliary metadata such as review topics, rating levels, text length, sentence position, app size, installation counts, and reviewer numbers (Guzman et al., 2015; Maalej and Nabil, 2015; Peng et al., 2016; Arya et al., 2019; Aslam et al., 2020; Merten et al., 2016a; Tizard et al., 2019). Additional auxiliary features mined via NLP include keywords (Maalej and Nabil, 2015; Peng et al., 2016; Song et al., 2020; Merten et al., 2016a), sentiment analysis (Williams and Mahmoud, 2017; Guzman et al., 2015; Maalej and Nabil, 2015; Peng et al., 2016; Aslam et al., 2020; Yadav et al., 2020; Panichella et al., 2015, 2016), semantic rules (Maalej and Nabil, 2015; Song et al., 2020), heuristic sentence identification (Panichella et al., 2015, 2016; Sorbo et al., 2015), lexical features (Panichella et al., 2016), and part-of-speech (POS) tagging (Panichella et al., 2015, 2016; Sorbo et al., 2015).
@@ 5.1.4. Pre-trained Language Model Approaches
Despite their effectiveness in handling large-scale data, PLMs exhibit a black-box nature, making their decision processes opaque. This opacity hinders developers’ ability to fully trust and utilize model outputs, especially considering potential biases in the training data. To address these issues, recent research has emphasized the importance of interpretability and transparency. For instance, Gambo et al. (Gambo et al., 2024) proposed a method that combines a BiLSTM with attention mechanisms and leverages local interpretable model-agnostic explanations (LIME) and SHapley Additive exPlanations (SHAP) to improve feature request identification. This approach outperforms SVM and CNN in accuracy and F1 score. Additionally, Biswas et al. (Biswas et al., 2024) developed the IARCT method, which quantifies the contribution of each word via Shapley values and provides contextual explanations for Transformer-based classification decisions.
@@ 5.1.5. General Large Language Model Approaches
Ghosh et al. (Ghosh et al., 2024) evaluated the performance of two PLMs (BERT and DistilBERT) and one LLM (GEMMA) on 3,200 Google Play Store user reviews. Their study confirmed the effectiveness of LLMs in parsing unstructured text and accurately identifying feature requests, showcasing the potential of LLMs to transform user feedback into actionable requirements. Building on this, Assi et al. (Assi et al., 2024) introduced the LLM-Cure framework, which derives feature improvement suggestions from user reviews via a two-phase approach. First, it extracts high-frequency functional features by semantically clustering batch reviews and normalizing cross-app features using word embeddings and cosine similarity. Then, for low-rated reviews of a target app, it retrieves corresponding high-rated competitor reviews to create contextual input for generating specific improvement suggestions through retrieval-augmented generation. Addressing challenges such as labor-intensive manual analysis and low clustering accuracy, Ren et al. (Ren et al., 2024) proposed a hybrid method that combines Latent Dirichlet Allocation (LDA) topic modeling for user review classification with a GPT-4–based generation approach for categories with sparse reviews. Furthermore, Wei et al. (Wei et al., 2024) conducted a comparative study of feature refinement approaches based on app store data and LLMs. Selecting 20 high-level root features (half existing and half novel) across diverse domains, they utilized App Store search and GPT-4 to generate two-level feature trees. Their integrated approach, implemented in the piStar (Pimentel and Castro, 2018) tool, reduces effort and inspires creativity, providing practical guidance for selecting appropriate feature request identification methods.
@@ 5.1.6. Other Approaches
Rahimi et al. (Rahimi and Cleland-Huang, 2014) proposed a semi-automated method for identifying feature requests from online forums by clustering user comments via incremental diffusion clustering and applying Apriori association rule mining to identify stakeholder-relevant topic clusters. Zhang et al. (Zhang et al., 2019b) developed a hybrid similarity-based approach to classify issue reports as bugs, feature requests, or others by combining normalized text similarities between unlabeled issue reports, user reviews, and labeled reports with weighted contributions. Shi et al. (Shi et al., 2020) introduced FRMiner, which uses a deep Siamese network to determine whether chat messages belong to the same class by learning their similarity, enabling feature requests identification. Panichella et al. (Panichella et al., 2015) combined manual rule matching of 246 linguistic patterns with sentiment analysis and J48 classification to categorize user reviews into information provision, search, feature requests, problem detection, and others. Gribkov et al. (Gribkov and Yekhlakov, 2020) presented the Trans-LSTM model to extract (topic, description) expressions from user review sentences, classifying them into bug reports, feature requests, and positive or negative features. Their use of Trans-CNN and Trans-LSTM outperformed prior models. Shi et al. (Shi et al., 2021) proposed an automated method for identifying feature requests from development emails using semantic sequence patterns generated from fuzzy rules and email context represented via a 2-gram model, validated on Ubuntu community emails. Izadi et al. (Izadi et al., 2022) introduced a multi-task approach that predicts both issue type (bug, feature, question) using fine-tuned BERT, compared with RF and SVM, and issue priority (high, medium, low), incorporating the predicted issue type to enhance cross-task learning.
Feature identification focuses on transforming unstructured user feedback into explicit and actionable feature requests that guide subsequent localization and implementation. Nadeem et al. (Nadeem et al., 2021) introduced a labeled dataset and evaluated supervised learning models, including LR, DT, and SVM, to identify software change request sentences in app reviews, demonstrating high effectiveness across models. Wu et al. (Wu et al., 2021) proposed KEFE, which extracts feature-describing phrases from app descriptions, aligns them with semantically related user reviews using a deep classifier, and applies regression analysis to measure their correlation with app ratings. AI-Safoury et al. (Al-Safoury et al., 2022) integrated app-store reviews with Github issue reports into unified change request lists through syntactic and sentiment-based classification, similarity-driven deduplication, and prioritization. Finally, Cai et al. (Cai et al., 2025) provided a systematic review of user-generated-content-based requirements elicitation, outlining how rule-based, topic modeling, and machine learning techniques have converged into a data-driven framework for automated feature identification.
@@ 5.2. Quality Review
Existing research on feature request review primarily centers on defining quality standards (Heck and Zaidman, 2015; Shi et al., 2016) and employing these criteria to assess the quality of requests (Heck and Zaidman, 2017, 2013). In addition, a growing body of work has proposed automated techniques to detect duplicate (Aversano, 2019; Gu et al., 2011; Li et al., 2017; Heck and Zaidman, 2014), redundant (Shi et al., 2016), and ambiguous (Aslam et al., 2020) feature requests, aiming to streamline the review process and improve the overall quality of requirements.